OpenStudy (anonymous):

A Pentagon has an interior that sum to_____ 360 540 720 900

OpenStudy (anonymous):

Pentagon has how many sides??

OpenStudy (anonymous):

if you know that will become your n.. Then use: \[Interior \; Angle = (n-2) \cdot 180^{\circ}\]

OpenStudy (anonymous):

(n-2)*180 so in this case 5-2=3*180=540

OpenStudy (anonymous):

thank u both

OpenStudy (anonymous):

you are welcome dear...

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

thanks :)